OpenStudy (diamond77108):

Plz help −(−y+18)+11=1

OpenStudy (mangotangochick):

First expand the brackets: -(−y+18)+11=1 goes to y-18+11=1 which is: y-7=1 Then get y on its own by adding +7 to both sides: y-7+7=1+7 which then goes to y=8 :)

alones (alones):

Yes \( y=8\) would be right.